# thanks: Corresponding Author

### v1

@@ Abstract
Axial-vector hadronic molecules $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast-}$ with the quark content $bb\overline{b}\overline{c}$ are studied using QCD sum rule method. The spectroscopic parameters of these molecules are computed in the context of the two-point sum rule method. Predictions for their masses are identical to each other and confirm that they are structures unstable against dissociations to ordinary heavy mesons. We evaluate the width of the state $\mathcal{M}_{\mathrm{AV}}$ and assume that it is equal to that of $\widetilde{\mathcal{M}}_{\mathrm{AV}}$. To this end, we explore its dominant decay channels $\mathcal{M}_{\mathrm{AV}}\to\Upsilon B_{c}^{-}$ and $\mathcal{M}_{\mathrm{AV}}\to\eta_{b}B_{c}^{\ast-}$. There also are subleading modes of $\mathcal{M}_{\mathrm{AV}}$ generated due to annihilation of $\overline{b}b$ quarks. We consider decays of the molecule $\mathcal{M}_{\mathrm{AV}}$ to pairs of the mesons $B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$. To find strong couplings at the $\mathcal{M}_{\mathrm{AV}}$-meson-meson vertices which determine the partial widths of these processes, we apply QCD three-point sum rule approach. The mass $m=(15800\pm 90)~\mathrm{MeV}$ and width $\Gamma[\mathcal{M}_{\mathrm{AV}}]=(112\pm 17)~\mathrm{MeV}$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ are useful for experimental studies of fully heavy molecular structures at ongoing and planning experiments.
@@ I Introduction
Hadronic molecules with the nonsymmetrical contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ establish another group of fully heavy particles. The molecules with these contents also attracted interest of researches Liu:2024pio ; Agaev:2025did . In fact, in Ref. Agaev:2025did we calculated the mass and decay width of the scalar states $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$: Our results demonstrated that they are relatively broad compounds.
@@
The correlation function contains two Lorentz structures. Because the term $\sim g_{\mu\nu}$ receives contributions arising only from the spin-1 particle, we choose it to carry our the sum rule analysis. Then the factor $\Lambda^{2}/(m^{2}-p^{2})$ is the invariant amplitude $\Pi^{\mathrm{Phys}}(p^{2})$ necessary for future studies.
@@
Here, $\rho^{\mathrm{OPE}}(s)$ is the spectral density found as an imaginary part of the function $\Pi^{\mathrm{OPE}}(p^{2})$. In the current work we take into account contributions to $\Pi^{\mathrm{OPE}}(p^{2})$ coming from the perturbative and dimension-four $\sim\langle\alpha_{s}G^{2}/\pi\rangle$ terms , therefore $\rho^{\mathrm{OPE}}(s)$ consists of $\rho^{\mathrm{pert.}}(s)$ and $\rho^{\mathrm{Dim4}}(s)$ components. The function $\Pi(M^{2})$ is extracted from the correlation function $\Pi^{\mathrm{OPE}}(p)$ and does not contain terms included into $\rho^{\mathrm{OPE}}(s)$.
@@ III.1 Decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$
The function $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ is a sum of two different Lorentz terms, one of which has be chosen for investigations. We continue with the invariant amplitude $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ that corresponds to the structure $g_{\mu\nu}$.
@@
Figure 4: QCD data and extrapolating functions $\mathcal{F}_{1}(Q^{2})$ (solid curve) and $\mathcal{F}_{2}(Q^{2})$ (dashed line). The diamond and circle fix the points $Q^{2}=-m_{B_{c}}^{2}$ and $Q^{2}=-m_{B_{c}^{\ast}}^{2}$, respectively.
@@ III.2 Process $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$
$\mathcal{F}_{2}^{0}=0.433~\mathrm{GeV}^{-1},l_{2}^{1}=4.47,\text{and }l_{2}^{2}=-9.31.$ (39)
The SR data and the function $\mathcal{F}_{2}(Q^{2})$ are plotted in Fig. 4 as well. We obtain for $g_{2}$
$g_{2}\equiv\mathcal{F}_{2}(-m_{B_{c}}^{2})=(1.7\pm 0.3)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (40)
@@
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}\right]=(31.7\pm 9.5)~\mathrm{MeV}.$ (42)
@@ IV.1 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$
In Eq. (45) $m_{B^{\ast}}=(5324.75\pm 0.20)~\mathrm{MeV},\ m_{\overline{D}^{0}}=(1864.84\pm 0.05)~\mathrm{MeV}$ and $f_{B^{\ast}}=(210\pm 6)~\mathrm{MeV},$ $f_{D}=(211.9\pm 1.1)~\mathrm{MeV}$ are the masses and decay constants of the these mesons, whereas $\varepsilon_{\mu}$ is the polarization vector of the vector meson $B^{\ast-}$. The vertex $\langle B^{\ast-}(p^{\prime},\varepsilon)\overline{D}^{0}(q)|B^{\ast-}(p,\epsilon)\rangle$ and correlator $\overline{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ are analogous to ones presented in section III.
@@
Results for $G(Q^{2})$ are plotted in Fig. 5. The extrapolating function $\overline{\mathcal{F}}(Q^{2})$ is determined by the constants $\overline{\mathcal{F}}^{0}=0.0166~\mathrm{GeV}^{-1}$, $\overline{l}^{1}=13.28$, and $\overline{l}^{2}=-16.70.$ The strong coupling $G$ is evaluated at the mass shell $q^{2}=m_{\overline{D}^{0}}^{2}$ and equal to
@@ IV.2 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$
These channels are studied within the same approach. Here, we give formulas for the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$ and our estimate its partial width: The partial width of the second mode $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{0}D^{\ast-}$ is approximately equal to parameters of the first decay.
@@
$\langle 0|J_{\mu}^{B^{\ast}}|\overline{B}_{s}^{\ast 0}(p^{\prime},\varepsilon)\rangle=f_{B_{s}^{\ast}}m_{B_{s}^{\ast}}\varepsilon_{\mu},$
@@
The vertex $\langle\overline{B}^{\ast 0}(p^{\prime},\varepsilon)D_{s}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle$ has the standard form. In Eq. (61) $m_{B_{s}^{\ast}}=(5415.4\pm 1.4)~\mathrm{MeV}$, $f_{B_{s}^{\ast}}=221~\mathrm{MeV}$ and $m_{D_{s}}=(1968.35\pm 0.07)~\mathrm{MeV}$, $f_{D_{s}}=(249.9\pm 0.5)~\mathrm{MeV}$ are the parameters of the final-state mesons.
@@
The SR for the form factor $\widehat{G}_{1}(Q^{2})$ is found by utilizing amplitudes connected with the terms $\sim g_{\mu\nu}$ both in $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ and $\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$.
@@
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\right]=(112\pm 17)~\mathrm{MeV}.$ (69)
@@ V Analysis and Final Notes
Investigations performed in the present work provide interesting information on parameters of the fully heavy axial-vector hadronic molecules, but, at the same time, give rise to different questions which should be clarified in the context of presumably alternative models and methods.

### v2

# Axial-vector molecules $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$

@@ Abstract
Axial-vector hadronic molecules $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast-}$ with the quark content $bb\overline{b}\overline{c}$ are studied using QCD sum rule method. The spectroscopic parameters of these molecules are computed in the context of the two-point sum rule method. Predictions for their masses are identical to each other and confirm that they are structures unstable against dissociations to ordinary heavy mesons. We evaluate the width of the state $\mathcal{M}_{\mathrm{AV}}$ and assume that it is equal to that of $\widetilde{\mathcal{M}}_{\mathrm{AV}}$. To this end, we explore its dominant decay channels $\mathcal{M}_{\mathrm{AV}}\to\Upsilon B_{c}^{-}$ and $\mathcal{M}_{\mathrm{AV}}\to\eta_{b}B_{c}^{\ast-}$. There also are subleading modes of $\mathcal{M}_{\mathrm{AV}}$ generated due to annihilation of $\overline{b}b$ quarks. We consider decays of the molecule $\mathcal{M}_{\mathrm{AV}}$ to pairs of the mesons $B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$. To find strong couplings at the $\mathcal{M}_{\mathrm{AV}}$-meson-meson vertices which determine the partial widths of these processes, we apply QCD three-point sum rule approach. The mass $m=(15800\pm 90)~\mathrm{MeV}$ and width $\Gamma[\mathcal{M}_{\mathrm{AV}}]=(114\pm 17)~\mathrm{MeV}$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ are useful for experimental studies of fully heavy molecular structures at ongoing and planning experiments.
@@ I Introduction
Hadronic molecules with the nonsymmetrical contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ establish another group of fully heavy particles. It is worth noting that, information about these structures is rather limited. Thus, these states are yet not observed in experiments. There are a few publications devoted to theoretical analyses of such molecules Liu:2024pio ; Agaev:2025did . In our work Agaev:2025did we calculated the mass and decay width of the scalar states $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$: Our results demonstrated that they are relatively broad compounds. But at lower limit of its mass $\mathcal{M}_{\mathrm{b}}$ may form a bound state of mesons $\eta_{b}$ and $B_{c}^{-}$. These interesting facts make the molecules with the contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ attractive objects for researches.
@@
The correlation function contains two Lorentz structures. Because the term $\sim g_{\mu\nu}$ receives contributions from the spin-1 particle, we choose it to carry our the sum rule analysis. Then the factor $\Lambda^{2}/(m^{2}-p^{2})$ is the invariant amplitude $\Pi^{\mathrm{Phys}}(p^{2})$ necessary for future studies.
@@
Here, $\rho^{\mathrm{OPE}}(s)$ is the spectral density found as an imaginary part of the function $\Pi^{\mathrm{OPE}}(p^{2})$. In the current work we take into account contributions to $\Pi^{\mathrm{OPE}}(p^{2})$ coming from the perturbative and dimension-four terms $\sim\langle\alpha_{s}G^{2}/\pi\rangle$. The reason is that dimension-$6$ contributions which are proportional to the triple-gluon condensate $\langle g_{s}^{3}G^{3}\rangle$ in the case of the heavy hadronic molecules are negligibly small: This was demonstrated by explicit calculations in Ref. Agaev:2025did when exploring the scalar molecule $\eta_{b}B_{c}^{-}$. Therefore, we truncate $\mathrm{OPE}$ at $\langle\alpha_{s}G^{2}/\pi\rangle$ level and neglect higher-dimension terms. As a result, $\rho^{\mathrm{OPE}}(s)$ consists of $\rho^{\mathrm{pert.}}(s)$ and $\rho^{\mathrm{Dim4}}(s)$ components. The function $\Pi(M^{2})$ in Eq. (10) is extracted from the correlation function $\Pi^{\mathrm{OPE}}(p)$ and does not contain terms included into $\rho^{\mathrm{OPE}}(s)$.
@@ III.1 Decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$
The function $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ is a sum of different Lorentz terms, one of which has be chosen for investigations. We continue with the invariant amplitude $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ that corresponds to the structure $g_{\mu\nu}$.
@@
Figure 4: QCD data and extrapolating function $\mathcal{F}_{1}(Q^{2})$. The diamond fixes the point $Q^{2}=-m_{B_{c}}^{2}$.
@@ III.2 Process $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$
$\mathcal{F}_{2}^{0}=0.276~\mathrm{GeV}^{-1},l_{2}^{1}=2.793,\text{and }l_{2}^{2}=-1.384.$ (39)
The SR data and the function $\mathcal{F}_{2}(Q^{2})$ are plotted in Fig. 5. We obtain for $g_{2}$
$g_{2}\equiv\mathcal{F}_{2}(-m_{B_{c}}^{2})=(1.69\pm 0.30)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (40)
@@
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}\right]=(33.3\pm 9.5)~\mathrm{MeV}.$ (42)
While considering the decay channels $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ we have applied the function Eq. (24) to extrapolate SR data to region of $Q^{2}<0$ which is necessary for estimation of the strong couplings at the relevant three-particle vertices. But this procedure can be fulfilled using alternative fit functions: A choice of a new function, in general, may modify results. To explore this problem, we introduce the function
$\mathcal{F}_{2A}(Q^{2})=\frac{f_{0}\left(1-Q^{2}/m^{2}\right)^{-2}}{\left[1-s_{1}(Q^{2}/m^{2})+s_{2}\left(Q^{4}/m^{4}\right)\right]},$ (43)
and compute the strong coupling $g_{2}$. Here, $f_{0}$, $s_{1}$ and $s_{2}$ are fitting constants. By employing SR data and Eq. (43) one gets $f_{0}=0.276~\mathrm{GeV}^{-1}$, $s_{1}=0.775$ and $s_{2}=2.486$. In Fig. 5 we demonstrate $\mathcal{F}_{2A}(Q^{2})$, where its nice agreement with SQ data, as well as with $\mathcal{F}_{2}(Q^{2})$ is evident. The extrapolating function $\mathcal{F}_{2A}(Q^{2})$ leads to the following prediction for the strong coupling $g_{2}=0.172~\mathrm{GeV}^{-1}$. The deviation $|0.003|$ of this value from one presented in Eq. (40) is an order of magnitude smaller that uncertainties $\pm 0.03$ of $g_{2}$. Therefore, throughout this article we utilize Eq. (24) and neglect small ambiguities connected with a choice of other extrapolating functions.
Figure 5: SR results and fitting functions $\mathcal{F}_{2}(Q^{2})$ (solid line) and $\mathcal{F}_{2A}(Q^{2})$ (dashed line). The star and triangle are at $Q^{2}=-m_{B_{c}^{\ast}}^{2}$.
@@ IV.1 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$
In Eq. (46) $m_{B^{\ast}}=(5324.75\pm 0.20)~\mathrm{MeV},\ m_{\overline{D}^{0}}=(1864.84\pm 0.05)~\mathrm{MeV}$ and $f_{B^{\ast}}=(210\pm 6)~\mathrm{MeV},$ $f_{D}=(211.9\pm 1.1)~\mathrm{MeV}$ are the masses and decay constants of these mesons, whereas $\varepsilon_{\mu}$ is the polarization vector of the vector meson $B^{\ast-}$. The vertex $\langle B^{\ast-}(p^{\prime},\varepsilon)\overline{D}^{0}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle$ and correlator $\overline{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ are analogous to ones presented in section III.
@@
Results for $G(Q^{2})$ are plotted in Fig. 6. The extrapolating function $\overline{\mathcal{F}}(Q^{2})$ is determined by the constants $\overline{\mathcal{F}}^{0}=0.0166~\mathrm{GeV}^{-1}$, $\overline{l}^{1}=13.28$, and $\overline{l}^{2}=-16.70.$ The strong coupling $G$ is evaluated at the mass shell $q^{2}=m_{\overline{D}^{0}}^{2}$ and equals to
@@ IV.2 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$
These channels are studied within the same approach. Here, we give formulas for the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$ and our estimate for its partial width: The partial width of the second mode $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{0}D^{\ast-}$ is approximately equal to parameters of the first decay.
@@
$\langle 0|J_{\mu}^{B^{\ast}_{s}}|\overline{B}_{s}^{\ast 0}(p^{\prime},\varepsilon)\rangle=f_{B_{s}^{\ast}}m_{B_{s}^{\ast}}\varepsilon_{\mu},$
@@
The vertex $\langle\overline{B}_{s}^{\ast 0}(p^{\prime},\varepsilon)D_{s}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle$ has the standard form. In Eq. (62) $m_{B_{s}^{\ast}}=(5415.4\pm 1.4)~\mathrm{MeV}$, $f_{B_{s}^{\ast}}=221~\mathrm{MeV}$ and $m_{D_{s}}=(1968.35\pm 0.07)~\mathrm{MeV}$, $f_{D_{s}}=(249.9\pm 0.5)~\mathrm{MeV}$ are the parameters of the final-state mesons.
@@
The SR for the form factor $\widehat{G}_{1}(Q^{2})$ is found by utilizing amplitudes connected with the terms $\sim g_{\mu\nu}$ both in $\widehat{\Pi}_{\mu\nu}^{\mathrm{1Phys}}(p,p^{\prime})$ and $\widehat{\Pi}_{\mu\nu}^{\mathrm{1OPE}}(p,p^{\prime})$.
@@
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\right]=(114\pm 17)~\mathrm{MeV}.$ (70)
@@ V Analysis and Final Notes
Studies performed in present article has important consequences for understanding features of the hadronic molecules $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$. It has been demonstrated that heavy mesons $\Upsilon$ and $B_{c}^{-}$ (as well as, $\eta_{b}$ and $B_{c}^{\ast-}$) may form a bound molecular state which does not dissociate to constituent particles. But due to annihilation of $b\overline{b}$ quarks it is unstable structure and has relatively small full width. Alternatively, the mesons $\Upsilon$ and $B_{c}^{-}$ may establish a broad compound. The first scenario is consistent with findings of Ref. Liu:2024pio . Evidently, these theoretical problems should be confirmed in the context of presumably alternative models and methods.
Obtained results are also important for experimental investigations of fully heavy structures. They provide information about masses and possible decay channels of hadronic molecules $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ pointing out final-states where they may be discovered. Thus, observation of an enhancement (a peak) in the mass distributions of the meson pairs $B^{\ast-}\overline{D}^{0}$, $B^{-}\overline{D}^{\ast 0}$ and others from Table 1 would be considered as evidence for realizations of the first scenario for $\mathcal{M}_{\mathrm{AV}}$, whereas the similar effect for mesons $\Upsilon B_{c}^{-}$ interpreted as correctness of the second option. Production mechanisms of heavy molecules $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ are interesting problems as well. But their investigations require independent and detailed investigations which are beyond the scope of the current work.